OpenStudy (wondermath):

How do you find the zeroes of cube root (x^2-1)

OpenStudy (anonymous):

\[\sqrt[3]{x^2-1}=\sqrt[3]{(x-1)(x+1)}=0\] cube both sides if you feel like it, not really necessary. \[(x-1)(x+1)=0, x=1, -1\]